OpenStudy (highschoolmom2010):

using Eulers formula find the missing value. F+V=E+2 faces:8 vertices:6 edges: ?

OpenStudy (anonymous):

F + V = E +2, you just have to plug in the given numbers aces:8 vertices:6 8 + 6 = E + 2 then solve for E (8+6) = 14 14= E + 2 (-2) you solve the rest... 14-2= E! :D

OpenStudy (highschoolmom2010):

i did that and it said the answer was not 12

OpenStudy (anonymous):

Really? thats weird

OpenStudy (anonymous):

then I have no idea... I am sorry :( maybe it is a computer issue

OpenStudy (highschoolmom2010):

said it was 8

OpenStudy (anonymous):

:o i think there is a computer issue with your software or something. Do you have a teacher ? you can ask them if it is a typo

OpenStudy (highschoolmom2010):

i just emailed her bc 14-2 DOESNT equal 8

OpenStudy (anonymous):

hahah! no it does not! lol awesome

OpenStudy (highschoolmom2010):

i am keeping it as 12